OpenStudy (anonymous):

There are 6 boys and 2 girls in a room. The mean of all their ages is 15. The mean of the girl's ages is 16.5. Work out the mean of the boys ages

OpenStudy (blurbendy):

Two equations (6B + 2G) / 8 = 15 GIVEN 8 * 15 = 120 ^ 8 because 6 boys + 2 girls = 8 and 15 because the mean is 15 2G/2 = 16.5 G = 16.5 GIVEN 6B + 2 * 16.5 = 120 6B + 33 = 120 6B = 87 B = 14.5

OpenStudy (anonymous):

what does 6B+33 mean @blurbendy

OpenStudy (blurbendy):

B is a variable. we dont know the mean age of the boys, so that is what B is trying to find. 33 because we know G = 16.5 2 girls * 16.5 = 33

OpenStudy (anonymous):

whats the total of boys ages looking at this @blurbendy

OpenStudy (blurbendy):

14.5 * 6